Parth (parthkohli):

How to determine what you should substitute for \(u\) and \(du\)? (Substitution Rule)

OpenStudy (lgbasallote):

you don't substitute anything to du...

OpenStudy (lgbasallote):

du is derivative of u

Parth (parthkohli):

Let me start with an example:\[ \int\left(1 - {1 \over w} \right) \cos (w -\ln w)dw\]

Parth (parthkohli):

Oh, yeah.

OpenStudy (lgbasallote):

you can start by letting u = 1/w

Parth (parthkohli):

How do you know what to substitute for \(u\)?

OpenStudy (lgbasallote):

if you're used to derivatives...then you can see them...

OpenStudy (lgbasallote):

on second thought u= 1/w is wrong

OpenStudy (lgbasallote):

let u = w - ln w

Parth (parthkohli):

\[\int f(g(x))g'(x) dx= \int f(u)du \text{ where }u = g(x) \]

Parth (parthkohli):

So in this case, \(\cos w = f(w)\) and \(w - \ln w = g(w)\)

Parth (parthkohli):

And the rule says that \(u = g(w)\) so \(u = w - \ln w\), right?

OpenStudy (lgbasallote):

yes u = w - ln w

Parth (parthkohli):

OK, thank you very much. :)